OpenStudy (anonymous):

Three forces act upon an object, which remains at rest. The forces are represented by the vectors P, Q and R, where P = 5i +4j and Q = -3i +6j Calculate the magnitude and direction in degree of vector R. I have done this and got magnitude 10.2 and direction 78.7 degree's Could someone please verify this or let me know where I have gone wrong. Thanks.

OpenStudy (anonymous):

Add PandQ, R and PQ must be of same magnitude and opposite angles

OpenStudy (anonymous):

Sorry Adrian I don't follow what you are saying. I have added P and Q to get R and made the magnitute 10.2.

OpenStudy (anonymous):

When you add up three forces, you net forces should be zero P = 5i +4j and Q = -3i +6j P={5,4} Q={-3,6} R={-2,-10}

OpenStudy (anonymous):

find mag and direction of R vector

OpenStudy (anonymous):

http://www.scribblar.com/kxnfusdm3

OpenStudy (anonymous):

\[\sqrt{4+100}=\sqrt{104}\] =\[2\sqrt{26}\]

OpenStudy (anonymous):

I am happy with the magnitude of 10.2 but I am being told that the direction is wrong. can you verify this??

OpenStudy (anonymous):

2 sqrt[26] evaluate to be 10.2 \[arctan(10/2)\]=78.69 but that's wrong , remember unit circle 180-78.69=101.31

OpenStudy (anonymous):

I dont have an option for 101.31 I have options 78.7, -11.3, -101.3, 11.3, -78.7, -168.7. Surely the resultantvector will be between P and Q? So I don't understand where the need to take 180 - 78.7

OpenStudy (anonymous):

Oh, I saw thinking {-2,10} instead of {-2,-10} arctan(10/2)=78.7 but let's visualize it right this time let's visualize

OpenStudy (anonymous):

so 180 -78.7 =101.3 but we are doing negative angel so -101.3

OpenStudy (anonymous):

Ok, cool so the R vector is pulling in the opposit direction 'balancing' the forces of P and Q? Got it ... thanks a lot. Medal on its way.

OpenStudy (anonymous):

180+78.7 is also a right answer if you go other way

OpenStudy (anonymous):

Also remember : net force of zero doesn't mean a object is at rest but when item is at rest , net force is zero

OpenStudy (anonymous):

Yeah ... sorted. Thanks. I was ignoring that the R was pulling in opposing direction. Some reason my head was saying it was a resultant of P and Q. Not quite the same! :)